4. (5.A.9) Define T E L(F3) by T(21, 22, 3) (222,0, 523). Find all eigenvalues and eigenvectors of T
Answer #1

T(2,, Z2, Z3) - (222, 0,52₂) T(1,0,0) = (0.o.o) T(0,1,0) = (2,0,0) T(0,0,1) - (0,0,5) T= OP 07 To 5 Idil-o-) 1-d2o1 od o=0 Toeigenvalues are 05. [т-от] x-o Goo LEON 1 ООО WC ooe T-5T X=0 По-5 ООО. elgesvectors are -00 -oo